OpenStudy (anonymous):

if sin(3x+10)degrees = cos20degrees, find x

OpenStudy (jdoe0001):

http://ichthyosapiens.com/School/Math/Trig/ch05CofunctionIdentities.jpg <--- notice those

OpenStudy (ipwnbunnies):

Therefore, the cos(20) = sin(70). Do you see the connection? Solve for x.

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ sin(90^o-\square^o )=cos(\square^o )}} \\ \quad \\ \quad \\ sin[(3x-10)^o]=cos(20^o)\implies sin[(3x-10)^o]={\color{blue}{ sin(90^o-20^o)}} \\ \quad \\ sin[(3x-10)^o]={\color{blue}{ sin(70^o)}}\implies 3x-10=70\)